bob invested $20,00, part at 14% and part at 13%. If the total interest at the end of the year is $2,720, how much did he invest at 14%?
Answer by funmath(2933)

(interest)=(principle)(rate)(time) or i=prt, in this case t is 1, so i=pr.
The amount of principle (p)invested at 14%: x
Let the rest of the $20,000 be the principle (p)invested at 13%:20,000-x
rate (r) at 14%=14/100=.14
rate (r) at 13%=13/100=.13
Therefore the interest earned at 14% is: .14x
and the interest earned at 13% is: .13(20,000-x)
If you add the two interests together and you get:$2,720
So your equation is:
.14x+.13(20,000-x)=2,720 Sometimes it makes easier to eliminate the decimals by multiplying both sides of the equation by 100.
100(.14x+.13(20,000-x))=100(2,720)
100(.14)x+100(.13)(20,000-x)=100(2,720)
14x+13(20,000-x)=272,000
14x+13(20,000)-13(x)=272,000
14x+260,000-13x=272,000
(14-13)x+260,000=272,000
x+260,000-260,000=272,000-260,000
x=12,000
The amount of money invested at 14% is $12,000.